# VistaWise[Uncaptioned image]: Building Cost-Effective Agent with Cross-Modal Knowledge Graph for Minecraft

### v2

@@ Abstract
Honghao Fu1,2‡thanks: The work was done during an internship at HKUST(GZ)., Junlong Ren1‡, Qi Chai1, Deheng Ye3, Yujun Cai2, Hao Wang1thanks: Corresponding Author. Equal Contribution. 1The Hong Kong University of Science and Technology (Guangzhou) 2The University of Queensland, 3Tencent honghao.fu@uq.edu.au, {jren686, qchai315}@connect.hkust-gz.edu.cn dericye@tencent.com, yujun.cai@uq.edu.au, haowang@hkust-gz.edu.cn
@@ 1 Introduction
Building on this objective, several studies [13, 14, 15] propose the use of Multimodal LLMs (MLLMs) [16] for reasoning based on visual information, subsequently enhancing visual policies with the reasoning outcomes. In parallel, to mitigate performance degradation or hallucinations [17, 18, 19, 20] resulting from the LLM’s lack of domain-specific knowledge in virtual environments, finetuning the LLM with knowledge tailored to the target virtual world has become a widely adopted strategy [10, 15, 13, 11, 21]. However, several challenges exist: (1) The reliance on additional visual policies to predict actions diminishes the LLM’s decision-making potential. Moreover, this approach usually models visual input globally, incorporating irrelevant noise that results in unstable predictions for MNK operations [14]. (2) General visual policies or LLMs may underperform due to the lack of domain-specific knowledge. However, finetuning them with large-scale domain-specific data needs prohibitive development costs.
@@ 2.1 Agents in Minecraft
Agents based on LLMs. Compared to traditional visual models, LLMs excel in complex cognitive tasks and causal reasoning [27, 28, 29, 30, 31, 32, 33, 34, 35], driving the development of Minecraft agents based on LLM policies [9, 10]. To address limitations of LLMs in visual capabilities, Voyager [3] proposes to leverage environmental APIs to obtain precise data and execute high-level actions. With the emergence of MLLMs [36], subsequent works such as STEVE [11], LARM [37], MP5 [21], LLaMA-Rider [38], and Odyssey [1] further enhance agent performance by incorporating visual inputs and finetuning MLLMs with domain-specific data. However, these methods rely on certain APIs to gather information and execute actions, facing challenges in generalization, as not all environments provide such APIs.
@@ 4.4 Inference Costs
In addition to the costs associated with model training, managing inference overhead is crucial to enhancing the cost-efficiency of LLM-based agents. As illustrated in Figure 5, we quantify this overhead through the token consumption required to achieve various goals. For the ultimate goal of “obtain diamond,” omitting full entity attributes during external knowledge injection and disabling the MLLM’s visual capabilities leads to a reduction in token consumption by 30.6% and 41.0% respectively, while the agent performance shows no significant degradation as indicated in Tables 2 and 3. Moreover, by leveraging the lightweight graph-based representations of cross-modal information, our framework achieves budget-friendly inference. Based on GPT-4o pricing for 160 iterations, our framework incurs a cost of $\sim\$1.28$. In comparison, the earlier cost-transparent [^4] https://github.com/MineDojo/Voyager/blob/main/FAQ.md LLM policy-based agent Voyager, under the same conditions, costs around $\sim\$25$, achieving a 94.9% reduction in cost. These results validate the cost-effective nature of our method.

### v3

@@ Abstract
Honghao Fu1,2‡thanks: The work was done during an internship at HKUST(GZ)., Junlong Ren1‡, Qi Chai1, Deheng Ye3, Yujun Cai2, Hao Wang1thanks: Corresponding Author. Equal Contribution. 1The Hong Kong University of Science and Technology (Guangzhou) 2University of Queensland, 3Tencent honghao.fu@uq.edu.au, {jren686, qchai315}@connect.hkust-gz.edu.cn dericye@tencent.com, yujun.cai@uq.edu.au, haowang@hkust-gz.edu.cn
@@ 1 Introduction
Building on this objective, several studies [64, 5, 65] propose the use of Multimodal LLMs (MLLMs) [33, 73] for reasoning based on visual information, subsequently enhancing visual policies with the reasoning outcomes. In parallel, to mitigate performance degradation or hallucinations [38, 61, 62, 28] resulting from the LLM’s lack of domain-specific knowledge in virtual environments, finetuning the LLM with knowledge tailored to the target virtual world has become a widely adopted strategy [63, 65, 64, 77, 43]. However, several challenges exist: (1) The reliance on additional visual policies to predict actions diminishes the LLM’s decision-making potential. Moreover, this approach usually models visual [16, 15] input globally, incorporating irrelevant noise that results in unstable predictions for MNK operations [5]. (2) General visual policies or LLMs may underperform due to the lack of domain-specific knowledge. However, finetuning them with large-scale domain-specific data needs prohibitive development costs.
@@ 2.1 Agents in Minecraft
Agents based on LLMs. Compared to traditional visual models, LLMs excel in complex cognitive tasks and causal reasoning [40, 37, 47, 49, 48, 53, 75, 29, 30, 69, 70], driving the development of Minecraft agents based on LLM policies [78, 63]. To address limitations of LLMs in visual [17, 14, 41] capabilities, Voyager [55] proposes to leverage environmental APIs to obtain precise data and execute high-level actions. With the emergence of MLLMs [11, 71, 72, 74], subsequent works such as STEVE [77], LARM [31], MP5 [43], LLaMA-Rider [9], and Odyssey [34] further enhance agent performance by incorporating visual inputs and finetuning MLLMs with domain-specific data. However, these methods rely on certain APIs to gather information and execute actions, facing challenges in generalization, as not all environments provide such APIs.
@@ 4.4 Inference Costs
In addition to the costs associated with model training, managing inference overhead is crucial to enhancing the cost-efficiency of LLM-based agents. As illustrated in Figure 5, we quantify this overhead through the token consumption required to achieve various goals. For the ultimate goal of “obtain diamond,” omitting full entity attributes during external knowledge injection and disabling the MLLM’s visual capabilities leads to a reduction in token consumption by 30.6% and 41.0% respectively, while the agent performance shows no significant degradation as indicated in Tables 2 and 3. Moreover, by leveraging the lightweight graph-based representations of cross-modal information, our framework achieves budget-friendly inference. Based on GPT-4o pricing for 160 iterations, our framework incurs a cost of $\sim\mathdollar 1.28$. In comparison, the earlier cost-transparent [^4] https://github.com/MineDojo/Voyager/blob/main/FAQ.md LLM policy-based agent Voyager, under the same conditions, costs around $\sim\mathdollar 25$, achieving a 94.9% reduction in cost. These results validate the cost-effective nature of our method.